# The Rational Torsion Subgroup of $J_{0}(\mathfrak{p}^{r})$

### v1

@@ 2 The rational cuspidal divisor class group of $X_{0}(\mathfrak{n})$
In this section, we recall from [8]. Let $\mathfrak{n}\lhd A$ be a non-zero ideal. We also denote by $\mathfrak{n}$ the monic polynomial in $A$ generating the ideal. The rational cuspidal divisor class group $\mathcal{C}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$ is a subgroup of $J_{0}(\mathfrak{n})$ generated by the linear equivalence classes of the degree 00 rational cuspidal divisors on $X_{0}(\mathfrak{n})$. By Gekeler [5, (3.6)], the cusps of $X_{0}(\mathfrak{n})$ are in bijection with $\Gamma_{0}(\mathfrak{n})\backslash\mathbb{P}^{1}(K)$. Moreover, every cusp of $X_{0}(\mathfrak{n})$ has a representative $\begin{bmatrix}a\\ d\end{bmatrix}$ in $\Gamma_{0}(\mathfrak{n})\backslash\mathbb{P}^{1}(K)$, where $a,d\in A$ are monic, $d|\mathfrak{n}$, and $\gcd(a,\mathfrak{n})=1$. Such a cusp is called of height $d$, independent to the choice of $a$. Denote by $[0]:=\begin{bmatrix}1\\ 1\end{bmatrix}$ and $[\infty]:=\begin{bmatrix}1\\ \mathfrak{n}\end{bmatrix}$ the unique cusps of $X_{0}(\mathfrak{n})$ of height $1$ and $\mathfrak{n}$, respectively.
In the following, we fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 1$. The cusps of $X_{0}(\mathfrak{n})$ of the same height are conjugate over $K$ by [13, Lem. 3.1]. For a monic $d\mid\mathfrak{n}$, let $(P(\mathfrak{n})_{d})$ be the sum of all the cusps of $X_{0}(\mathfrak{n})$ of height $d$, which is a rational cuspidal divisor on $X_{0}(\mathfrak{n})$; cf. [8, Sec. 1.3]. Let
@@
be the group of the degree 00 rational cuspidal divisors on $X_{0}(\mathfrak{n})$. Recall that a modular function on $X_{0}(\mathfrak{n})$ is a meromorphic function on $\Omega\cup\ \mathbb{P}^{1}(K)$ which is invariant under $\Gamma_{0}(\mathfrak{n})$, and a modular unit on $X_{0}(\mathfrak{n})$ is a modular function on $X_{0}(\mathfrak{n})$ that does not have zeros or poles on $\Omega$. Let $\mathcal{U}_{\mathfrak{n}}$ be the subgroup of $\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$ consisting of the divisors of modular units. Let $C_{i}:=(P(\mathfrak{n})_{\mathfrak{p}^{i}})-\deg(P(\mathfrak{n})_{\mathfrak{p}^{i}})\cdot[\infty]\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$, where $10\leq i\leq r-$ 1. Then
@@
Let $\Delta(z)$ be the Drinfeld discriminant function defined in [5, p. 183]. For $d|\mathfrak{n}$, consider $\Delta_{d}(z):=\Delta(dz)$, which is a modular form on $\Omega$ of weight $q^{2}-1$ and type 00 for $\Gamma_{0}(\mathfrak{n})$; cf. [5, (1.2)]. The zero orders of $\Delta_{d}(z)$ at the cusps of $X_{0}(\mathfrak{n})$ are defined in [2, p. 47]. Using [5, Eq. (3.10) and (3.11)], we found the divisor of $\Delta_{d}(z)$ on $X_{0}(\mathfrak{n})$ in [8, Sec. 2.1]. Let $\mathcal{E}_{\mathfrak{n}}$ be the group of modular units on $X_{0}(\mathfrak{n})$. Recall the following group homomorphism in [8, Sec. 2.1]:
${g:}$${\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)}$$\tens{\displaystyle\mathcal{E}_{\mathfrak{n}}\tens_{\mathbb{Z}}\mathbb{Q}}$${C=\displaystyle\sum_{\begin{subarray}{c}d|\mathfrak{n}\\ \text{monic}\end{subarray}}a_{d}\cdot(P(\mathfrak{n})_{d})}$${\left(\displaystyle\prod_{\begin{subarray}{c}d|\mathfrak{n}\\ \text{monic}\end{subarray}}\Delta_{d}^{r_{d}}\right)\otimes\frac{1}{(q-1)(|\mathfrak{p}|^{2}-1)|\mathfrak{p}|^{r-1}},}$
@@
Here, $\Lambda(\mathfrak{n})^{-1}\in\operatorname{GL}_{r+1}(\mathbb{Q})$ is a certain matrix defined in [8, Sec. 2.1]. The map $g$ provides a connection between the rational cuspidal divisor class group $\mathcal{C}(\mathfrak{n})$ and the modular units on $X_{0}(\mathfrak{n})$. More precisely, if $g(C)=f\otimes a$ for some $C\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$, $f\in\mathcal{E}_{\mathfrak{n}}$, and $a\in\mathbb{Q}$, then we have $C=a\cdot\operatorname{div}(f)$.
@@ 3 Harmonic cochains and Fourier expansions
Let $\mathscr{T}$ be the Bruhat-Tits tree of $\operatorname{PGL}(2,K_{\infty})$; cf. [7, (1.3)]. Recall that $G$ is the group scheme $\operatorname{GL}(2)$ over $\mathbb{F}_{q}$, with $Z$ the group scheme of scalar matrices in $G$. Let $\mathcal{K}=G(\mathcal{O}_{\infty})$ with its Iwahori subgroup $\mathcal{I}$ defined by
@@
Let $V(\mathscr{T})=G(K_{\infty})/\mathcal{K}\cdot Z(K_{\infty})$ and $E(\mathscr{T})=G(K_{\infty})/\mathcal{I}\cdot Z(K_{\infty})$ be the sets of the vertices and the oriented edges of $\mathscr{T}$, respectively; cf. [7, (1.3)]. The group $G(K_{\infty})$ acts on $E(\mathscr{T})$ by left multiplication. For a function $f$ on $E(\mathscr{T})$ and $\gamma\in G(K_{\infty})$, we define a function $f|\gamma$ on $E(\mathscr{T})$ by $(f|\gamma)(e):=f(\gamma e)$.
@@
We introduce some notation. Fix a non-zero ideal $\mathfrak{n}\lhd A$. Denote by $\mathcal{H}(\mathscr{T},R)$ the group of $R$-valued harmonic cochains on $\mathscr{T}$ and by $\mathcal{H}(\mathfrak{n},R):=\mathcal{H}(\mathscr{T},R)^{\Gamma_{0}(\mathfrak{n})}$ its subgroup of $\Gamma_{0}(\mathfrak{n})$-invariant harmonic cochains, i.e. $f|\gamma=f$ for all $\gamma\in\Gamma_{0}(\mathfrak{n})$. For $f\in\mathcal{H}(\mathfrak{n},R)$, it naturally defines a function $f^{\prime}$ on the quotient graph $\Gamma_{0}(\mathfrak{n})\backslash\mathscr{T}$. If $f^{\prime}$ has compact support, then we call $f$ a cuspidal harmonic cochain for $\Gamma_{0}(\mathfrak{n})$. Let $\mathcal{H}_{0}(\mathfrak{n},R)$ be the subgroup of $\mathcal{H}(\mathfrak{n},R)$ consisting of cuspidal harmonic cochains for $\Gamma_{0}(\mathfrak{n})$, and denote by $\mathcal{H}_{00}(\mathfrak{n},R)$ the image of $\mathcal{H}_{0}(\mathfrak{n},\mathbb{Z})\otimes R$ in $\mathcal{H}_{0}(\mathfrak{n},R)$. Note that if $R$ is flat over $\mathbb{Z}$, then $\mathcal{H}_{0}(\mathfrak{n},R)=\mathcal{H}_{00}(\mathfrak{n},R)$.
Definition 3.2 (Pál [11, Sec. 2]).
A ring $R$ is called a coefficient ring if $1/p\in R$ and $R$ is a quotient of a discrete valuation ring $\widetilde{R}$ which contains $p$-th roots of unity.
@@
Example.
Recall that $\Delta\in\mathcal{O}(\Omega)^{\ast}$ is the Drinfeld discriminant function. Consider the function $\widetilde{r}(\Delta)\in\mathcal{H}(\mathscr{T},\mathbb{Z})\subset\mathcal{H}(\mathscr{T},\mathbb{C})$, which is invariant under $\Gamma_{\infty}$. By Gekeler [5, (2.5) and Cor. 2.8], we have
@@
Then we derive that:
@@
By the previous lemma, we have $\widetilde{r}(\Delta_{\mathfrak{p}^{i}})^{0}(\pi_{\infty}^{k})=\widetilde{r}(\Delta)^{0}(\pi_{\infty}^{k-\deg(\mathfrak{p}^{i})})$ and $\widetilde{r}(\Delta_{\mathfrak{p}^{i}})^{\ast}(1)=\widetilde{r}(\Delta)^{\ast}((\mathfrak{p}^{i})^{-1})$, which imply the result by Eq. (2). ∎
@@ 4 Hecke operators and Eisenstein ideals
Fix a prime-power ideal $\mathfrak{n}=\mathfrak{p}^{r}$ of $A$ with $r\geq 2$, and let $R$ be a commutative ring with unity. For a non-zero ideal $\mathfrak{m}\lhd A$, we define an $R$-linear transformation of the space of $R$-valued functions on $E(\mathscr{T})$ by
@@
where the sum is over $a,b,d\in A$ such that $a$, $d$ are monic, $(ad)=\mathfrak{m}$, $(a)+\mathfrak{n}=A$, and $\deg(b)<\deg(d)$. This transformation is the $\mathfrak{m}$-th Hecke operator; cf. [13, Sec. 2.3]. In particular, for the prime divisor $\mathfrak{p}$ of $\mathfrak{n}$, write
$f|U_{\mathfrak{p}}:=\sum_{\begin{subarray}{c}b\in A\\ \deg(b)<\deg(\mathfrak{p})\end{subarray}}f|\begin{pmatrix}1&b\\ 0&\mathfrak{p}\end{pmatrix},$
then we see that $U_{\mathfrak{p}}=T_{\mathfrak{p}}$.
@@
The definition of $T_{\mathfrak{p}}:=U_{\mathfrak{p}}$ here is different from that of $T_{\mathfrak{p}}:=U_{\mathfrak{p}}+B_{\mathfrak{p}}$ in [5, (1.17)].
Proposition 4.2 ([12, Prop. 2.10]).
@@
Example.
Consider the action of $U_{\mathfrak{p}}$ on $\mathcal{H}(\mathfrak{n},R)$. Let $f\in\mathcal{H}(\mathfrak{n},R)$. Since $\mathfrak{p}^{2}\mid\mathfrak{n}$, we have $h:=f|U_{\mathfrak{p}}\in\mathcal{H}(\mathfrak{n}/\mathfrak{p},R)$ by [13, Lem. 2.6]. Moreover, since $h$ is invariant under $\Gamma_{\infty}$, we have $(h|B_{\mathfrak{p}})|U_{\mathfrak{p}}=|\mathfrak{p}|\cdot h$ by [12, Lem. 2.23]. We obtain:
@@
Let $v$ be a place of $K$. Denote by $K_{v}$ the completion of $K$ at $v$, $\mathcal{O}_{v}$ the ring of integers of $K_{v}$, and $\mathbb{F}_{v}$ the residue field of $\mathcal{O}_{v}$. Moreover, we denote by $\mathbb{F}_{v}^{\text{nr}}$ and $\mathcal{O}_{v}^{\text{nr}}$ the maximal unramified extension of $\mathbb{F}_{v}$ and $\mathcal{O}_{v}$, respectively. Let $\mathcal{J}$ denote the Néron model of $J_{0}(\mathfrak{n})$ over $\mathcal{O}_{v}$. Let $\mathcal{J}^{0}$ denote the relative connected component of the identity of $\mathcal{J}$, that is, the largest open subscheme of $\mathcal{J}$ containing the identity section which has connected fibres. The group of connected components of $J_{0}(\mathfrak{n})$ at $v$ is $\Phi_{v}:=\mathcal{J}_{\mathbb{F}_{v}}/\mathcal{J}_{\mathbb{F}_{v}}^{0}$. There is a natural map in [12, p. 586], called the canonical specialization map:
$\rho_{v}:J_{0}(\mathfrak{n})(K_{v}^{\text{nr}})=\mathcal{J}(\mathcal{O}_{v}^{\text{nr}})\rightarrow\mathcal{J}_{\mathbb{F}_{v}}(\overline{\mathbb{F}_{v}})\rightarrow\Phi_{v}.$ (3)
The Hecke algebra $\mathbb{T}(\mathfrak{n})$ naturally acts on the Jacobian variety $J_{0}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$, which functorially extends to $\mathcal{J}$; hence, $\mathbb{T}(\mathfrak{n})$ also acts on $\Phi_{v}$; cf. [13, p. 386].

### v2

@@ 2 The rational cuspidal divisor class group of $X_{0}(\mathfrak{n})$
In this section, we recall from [9]. Let $\mathfrak{n}\lhd A$ be a non-zero ideal. We also denote by $\mathfrak{n}$ the monic polynomial in $A$ generating the ideal. The rational cuspidal divisor class group $\mathcal{C}(\mathfrak{n})$ of $X_{0}(\mathfrak{n})$ is the subgroup of $J_{0}(\mathfrak{n})$ generated by the linear equivalence classes of the degree 00 rational cuspidal divisors on $X_{0}(\mathfrak{n})$. By Gekeler [5, (3.6)], the cusps of $X_{0}(\mathfrak{n})$ are in bijection with $\Gamma_{0}(\mathfrak{n})\backslash\mathbb{P}^{1}(K)$. Moreover, every cusp of $X_{0}(\mathfrak{n})$ has a representative $\begin{bmatrix}a\\ d\end{bmatrix}$ in $\Gamma_{0}(\mathfrak{n})\backslash\mathbb{P}^{1}(K)$, where $a,d\in A$ are monic, $d|\mathfrak{n}$, and $\gcd(a,\mathfrak{n})=1$. Two representatives $\begin{bmatrix}a\\ d\end{bmatrix}$ and $\begin{bmatrix}a^{\prime}\\ d^{\prime}\end{bmatrix}$ represents the same cusp of $X_{0}(\mathfrak{n})$ if and only if $d=d^{\prime}$ and $\alpha a^{\prime}\equiv a\mod\gcd(d,\mathfrak{n}/d)$ for some $\alpha\in\mathbb{F}_{q}^{\times}$ by [14, Lemma 3.1 (i)]. Such a cusp is called of height $d$, independent of the choice of $a$. The cusps of $X_{0}(\mathfrak{n})$ of the same height are conjugate over $K$ by [7, Section 6] and [14, Lemma 3.1 (iii)]. Denote by $[0]$ and []delimited-[][$\infty$][ ∞ ] the unique cusp of $X_{0}(\mathfrak{n})$ of height $1$ and $\mathfrak{n}$, respectively.
In the following, we fix a prime power ideal $\mathfrak{n}=\mathfrak{p}^{r}\lhd A$ with $r\geq 1$. For a monic $d\mid\mathfrak{n}$, denote by $(P(\mathfrak{n})_{d})$ the sum of all the cusps of $X_{0}(\mathfrak{n})$ of height $d$, which is a rational cuspidal divisor on $X_{0}(\mathfrak{n})$; cf. [9, Section 1.3]. Let
@@
be the group of the degree 00 rational cuspidal divisors on $X_{0}(\mathfrak{n})$. Recall that a modular function on $X_{0}(\mathfrak{n})$ is a meromorphic function on $\Omega\cup\ \mathbb{P}^{1}(K)$ which is invariant under the action of $\Gamma_{0}(\mathfrak{n})$. Moreover, a modular unit on $X_{0}(\mathfrak{n})$ is a modular function on $X_{0}(\mathfrak{n})$ that does not have zeros or poles on $\Omega$. Let $\mathcal{U}_{\mathfrak{n}}$ be the subgroup of $\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$ consisting of the divisors of modular units. Let $C_{i}:=(P(\mathfrak{n})_{\mathfrak{p}^{i}})-\deg(P(\mathfrak{n})_{\mathfrak{p}^{i}})\cdot[\infty]\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$, where $10\leq i\leq r-$ 1. Then
@@
Let $\Delta(z)$ be the Drinfeld discriminant function defined in [5, p. 183]. For $d|\mathfrak{n}$, consider $\Delta_{d}(z):=\Delta(dz)$, which is a modular form on $\Omega$ of weight $q^{2}-1$ and type 00 for $\Gamma_{0}(\mathfrak{n})$; cf. [5, (1.2)]. The zero orders of $\Delta_{d}(z)$ at the cusps of $X_{0}(\mathfrak{n})$ are defined in [2, p. 47]. Using [5, Equations (3.10) and (3.11)], we found the divisor of $\Delta_{d}(z)$ on $X_{0}(\mathfrak{n})$ in [9, Section 2.1]. Let $\mathcal{O}(X_{0}(\mathfrak{n}))^{\ast}$ be the (multiplicative) group of modular units on $X_{0}(\mathfrak{n})$. Recall the following group homomorphism in [9, Section 2.1]:
${g:}$${\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)}$$\tens{\displaystyle\mathcal{O}(X_{0}(\mathfrak{n}))^{\ast}\tens_{\mathbb{Z}}\mathbb{Q}}$${C=\displaystyle\sum_{\begin{subarray}{c}d|\mathfrak{n}\\ \text{monic}\end{subarray}}a_{d}\cdot(P(\mathfrak{n})_{d})}$${\left(\displaystyle\prod_{\begin{subarray}{c}d|\mathfrak{n}\\ \text{monic}\end{subarray}}\Delta_{d}^{r_{d}}\right)\otimes\frac{1}{(q-1)(|\mathfrak{p}|^{2}-1)|\mathfrak{p}|^{r-1}},}$
@@
Here, $\Lambda(\mathfrak{n})^{-1}\in\operatorname{GL}_{r+1}(\mathbb{Q})$ is a certain matrix defined in [9, Section 2.1]. The map $g$ provides a connection between the rational cuspidal divisor class group $\mathcal{C}(\mathfrak{n})$ and the modular units on $X_{0}(\mathfrak{n})$. More precisely, if $g(C)=f\otimes a$ for some $C\in\operatorname{Div}_{\operatorname{cusp}}^{0}(X_{0}(\mathfrak{n}))(K)$, $f\in\mathcal{O}(X_{0}(\mathfrak{n}))^{\ast}$, and $a\in\mathbb{Q}$, then we have $C=a\cdot\operatorname{div}(f)$.
@@ 3 Harmonic cochains and Fourier expansions
Let $\mathscr{T}$ be the Bruhat-Tits tree of $\operatorname{PGL}(2,K_{\infty})$; cf. [8, (1.3)]. Recall that $G$ is the group scheme $\operatorname{GL}(2)$ over $\mathbb{F}_{q}$, with $Z$ the group scheme of scalar matrices in $G$. Let $\mathcal{K}=G(\mathcal{O}_{\infty})$ with its Iwahori subgroup $\mathcal{I}$ defined by
@@
Let $V(\mathscr{T})=G(K_{\infty})/\mathcal{K}\cdot Z(K_{\infty})$ and $E(\mathscr{T})=G(K_{\infty})/\mathcal{I}\cdot Z(K_{\infty})$ be the sets of the vertices and the oriented edges of $\mathscr{T}$, respectively; cf. [8, (1.3)]. The group $G(K_{\infty})$ acts on $E(\mathscr{T})$ by left multiplication. For a function $f$ on $E(\mathscr{T})$ and $\gamma\in G(K_{\infty})$, we define a function $f|\gamma$ on $E(\mathscr{T})$ by $(f|\gamma)(e):=f(\gamma e)$.
@@
We introduce some notation. Fix a non-zero ideal $\mathfrak{n}\lhd A$. Denote by $\mathcal{H}(\mathscr{T},R)$ the group of $R$-valued harmonic cochains on $\mathscr{T}$ and by $\mathcal{H}(\mathfrak{n},R):=\mathcal{H}(\mathscr{T},R)^{\Gamma_{0}(\mathfrak{n})}$ its subgroup of $\Gamma_{0}(\mathfrak{n})$-invariant harmonic cochains, i.e., $f|\gamma=f$ for all $\gamma\in\Gamma_{0}(\mathfrak{n})$. For $f\in\mathcal{H}(\mathfrak{n},R)$, it naturally defines a function $f^{\prime}$ on the quotient graph $\Gamma_{0}(\mathfrak{n})\backslash\mathscr{T}$. If $f^{\prime}$ has compact support, then we call $f$ a cuspidal harmonic cochain for $\Gamma_{0}(\mathfrak{n})$. Let $\mathcal{H}_{0}(\mathfrak{n},R)$ be the subgroup of $\mathcal{H}(\mathfrak{n},R)$ consisting of cuspidal harmonic cochains for $\Gamma_{0}(\mathfrak{n})$, and denote by $\mathcal{H}_{00}(\mathfrak{n},R)$ the image of $\mathcal{H}_{0}(\mathfrak{n},\mathbb{Z})\otimes R$ in $\mathcal{H}_{0}(\mathfrak{n},R)$. Note that if $R$ is flat over $\mathbb{Z}$, then $\mathcal{H}_{0}(\mathfrak{n},R)=\mathcal{H}_{00}(\mathfrak{n},R)$.
Definition 3.2 (Pál [12, Section 2]).
A ring $R$ is a coefficient ring if $1/p\in R$ and $R$ is a quotient of a discrete valuation ring $\widetilde{R}$ which contains $p$-th roots of unity.
@@
Example 1.
Recall that $\Delta\in\mathcal{O}(\Omega)^{\ast}$ is the Drinfeld discriminant function. Consider the function $\widetilde{r}(\Delta)\in\mathcal{H}(\mathscr{T},\mathbb{Z})\subset\mathcal{H}(\mathscr{T},\mathbb{C})$, which is invariant under $\Gamma_{\infty}$. By [5, (2.5) and Corollary 2.8], we have
@@
Note that $f^{\ast}(\mathfrak{n}/\mathfrak{m})=0$ if $\mathfrak{m}\nmid\mathfrak{n}$.
We derive that:
@@
By the previous lemma, we have $\widetilde{r}(\Delta_{\mathfrak{p}^{i}})^{0}(\pi_{\infty}^{k})=\widetilde{r}(\Delta)^{0}(\pi_{\infty}^{k-\deg(\mathfrak{p}^{i})})$ and $\widetilde{r}(\Delta_{\mathfrak{p}^{i}})^{\ast}(1)=\widetilde{r}(\Delta)^{\ast}((\mathfrak{p}^{i})^{-1})$, which imply the result by Equation (2). ∎
@@ 4 Hecke operators and Eisenstein ideals
Fix a prime power ideal $\mathfrak{n}=\mathfrak{p}^{r}$ of $A$ with $r\geq 2$, and let $R$ be a commutative ring with unity. For a non-zero ideal $\mathfrak{m}\lhd A$, we define an $R$-linear transformation of the space of $R$-valued functions on $E(\mathscr{T})$ by
@@
where the sum is over $a,b,d\in A$ such that $a$, $d$ are monic, $(ad)=\mathfrak{m}$, $(a)+\mathfrak{n}=A$, and $\deg(b)<\deg(d)$. This transformation is the $\mathfrak{m}$-th Hecke operator; cf. [14, Section 2.3]. In particular, for the prime divisor $\mathfrak{p}$ of $\mathfrak{n}$, we have
$f|U_{\mathfrak{p}}:=\sum_{\begin{subarray}{c}b\in A\\ \deg(b)<\deg(\mathfrak{p})\end{subarray}}f|\begin{pmatrix}1&b\\ 0&\mathfrak{p}\end{pmatrix}=f|T_{\mathfrak{p}}.$
@@
The definition of $T_{\mathfrak{p}}=U_{\mathfrak{p}}$ above is different from that of $T_{\mathfrak{p}}:=U_{\mathfrak{p}}+B_{\mathfrak{p}}$ in [5, (1.17)].
Proposition 4.2 ([13, Proposition 2.10]).
@@
Example 2.
Consider the action of $U_{\mathfrak{p}}$ on $\mathcal{H}(\mathfrak{n},R)$. Let $f\in\mathcal{H}(\mathfrak{n},R)$. Since $\mathfrak{p}^{2}\mid\mathfrak{n}$, we have $h:=f|U_{\mathfrak{p}}\in\mathcal{H}(\mathfrak{n}/\mathfrak{p},R)$ by [14, Lemma 2.6]. Moreover, since $h$ is invariant under $\Gamma_{\infty}$, we have $(h|B_{\mathfrak{p}})|U_{\mathfrak{p}}=|\mathfrak{p}|\cdot h$ by [13, Lemma 2.23]. We obtain:
@@
Let $v$ be a place of $K$. Denote by $K_{v}$ the completion of $K$ at $v$, $\mathcal{O}_{v}$ the ring of integers of $K_{v}$, $\mathbb{F}_{v}$ the residue field of $\mathcal{O}_{v}$, and $\overline{\mathbb{F}_{v}}$ an algebraic closure of $\mathbb{F}_{v}$. Moreover, we denote by $K_{v}^{\text{nr}}$ and $\mathcal{O}_{v}^{\text{nr}}$ the maximal unramified extension of $K_{v}$ and $\mathcal{O}_{v}$, respectively. Let $\mathcal{J}$ denote the Néron model of $J_{0}(\mathfrak{n})$ over $\mathbb{P}^{1}_{\mathbb{F}_{q}}$. Let $\mathcal{J}^{0}$ denote the relative connected component of the identity of $\mathcal{J}$, that is, the largest open subscheme of $\mathcal{J}$ containing the identity section which has connected fibres. The group of connected components of $J_{0}(\mathfrak{n})$ at $v$ is $\Phi_{v}:=\mathcal{J}_{\mathbb{F}_{v}}/\mathcal{J}_{\mathbb{F}_{v}}^{0}$. There is a natural map in [13, p. 586], called the canonical specialization map:
$\wp_{v}:J_{0}(\mathfrak{n})(K_{v}^{\text{nr}})=\mathcal{J}(\mathcal{O}_{v}^{\text{nr}})\rightarrow\mathcal{J}_{\mathbb{F}_{v}}(\overline{\mathbb{F}_{v}})\rightarrow\Phi_{v}.$ (3)
We remark that the Hecke algebra $\mathbb{T}(\mathfrak{n})$ naturally acts on $J_{0}(\mathfrak{n})$, which functorially extends to $\mathcal{J}$; hence, $\mathbb{T}(\mathfrak{n})$ also acts on $\Phi_{v}$; cf. [14, p. 386].